One True Love by Educational Level A study is conducted in which people were asked whether they agreed or disagreed with theRound your answer for the chi-square statistic to one decimal place, and your answer for the p-value to three decimal places.
Answer #1

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: Two categorical variables are independent.

Alternative hypothesis: Ha: Two categorical variables are dependent.

We assume level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 3

Number of columns = c = 3

Degrees of freedom = df = (r – 1)*(c – 1) = 2*2 = 4

α = 0.05

Critical value = 9.487729

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

HS

Some

College

Total

Agree

360

165

200

725

Disagree

557

467

787

1811

Don't know

15

28

28

71

Total

932

660

1015

2607

Expected Frequencies

Column variable

Row variable

HS

Some

College

Total

Agree

259.1868

183.5443

282.2689

725

Disagree

647.4308

458.481

705.0882

1811

Don't know

25.38243

17.97468

27.64288

71

Total

932

660

1015

2607

Calculations

(O - E)

100.8132

-18.5443

-82.2689

-90.4308

8.518987

81.91178

-10.3824

10.02532

0.357115

(O - E)^2/E

39.21226

1.873614

23.97774

12.63104

0.15829

9.515886

4.246831

5.591585

0.004614

Test Statistic = Chi square = ∑[(O – E)^2/E] = 97.21186

Chi-square statistic = 97.2

P-value = 0.000

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

Conclusion: Reject H0

Does the test indicate an association between these two variables?

Answer: Yes

The test indicates that there is an association between these two variables because we reject the null hypothesis.